OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 8, -14, and 3 + 9i

OpenStudy (anonymous):

first part is easy right? \[(x-8)(x+14)\] for the zeros of \(8\) and \(-14\) second part is annoying

OpenStudy (anonymous):

there is however a more or less snappy way to find the quadratic with the zeros \(3+9i\) and \(3-9i\)

OpenStudy (anonymous):

I know that you have to multiply the zeros together, but I don't know how to multiply (3+9i) and (3-9i) with the other zeros. could you help explain how to do that?

OpenStudy (anonymous):

one way is to work backwards start with \[x=3+9i\] then subtract 3 to get \[x-3=9i\] square both sides and get \[(x-3)^2=-81\] or \[x^2-6x+9=-81\] then add \(81\) and get \[x^2-6x+90\]

OpenStudy (anonymous):

@suzanna you do not multiply the zeros together

OpenStudy (anonymous):

you have to multiply all the factors together i.e. you have to multiply \[(x-8)(x+14)(x-(3+9i))(x-(3-9i))\] so as you can see it is kind of a drag

OpenStudy (anonymous):

to recap \((x-8)(x+14)\) is a routine multiplication right?

OpenStudy (anonymous):

yes I can see, and that is where I get confused as to how to multiply them together and yes that is easy!

OpenStudy (anonymous):

i will show you a really snappy way but again lets make it clear \((x-8)(x+14)\) is the easy part \((x-(3+9i)(x-(3-9i)\) is the hard part

OpenStudy (anonymous):

yes i got x^2+6x-112 for (x-8)(x=14)

OpenStudy (anonymous):

that is right , yes

OpenStudy (anonymous):

okay, now how would i do the rest? I just get confused after that.

OpenStudy (anonymous):

ok lets do it two ways you want the real fast way first? comes with no explanation, just mechanics

OpenStudy (anonymous):

yes please!

OpenStudy (anonymous):

suppose you want a quadratic polynomial with leading coefficient 1 and zero \(a+bi\) which of course means it also has the zero \(a-bi\) the polynomial is \[x^2-2ax+(a^2+b^2)\]

OpenStudy (anonymous):

for example, if you want a zero of \(4+3i\) then the quadratic is \[x^2-2\times 4x+(4^2+3^2)=x^2-8x+25\]

OpenStudy (anonymous):

in your case it will be \[x^2-2\times 3x+3^2+9^2=x^2-6x+90\]

OpenStudy (anonymous):

fast right?

OpenStudy (anonymous):

yes it is! Is that all you have to do?

OpenStudy (anonymous):

however, it comes at the disadvantage of having to memorize it and no explanation

OpenStudy (anonymous):

so let me repeat what i had in the first post suppose you want a zero of \(3+9i\) you can put \(x=3+9i\) and work backwards

OpenStudy (anonymous):

i.e. put \[x=3+9i\] subtract 3 and get \[x-3=9i\] square both sides to get \[(x-3)^2=-81\] or \[x^2-6x+9=-81\] then add \(81\) to both sides and get \[x^2-6x+90=0\] same thing obviously

OpenStudy (anonymous):

yes i see!

OpenStudy (anonymous):

of course if you have the stomach for it, you can always multiply \[(x-(3+9i)(x-(3-9i)\]which is not as bad as it looks first is \(x^2\) last is \((3+9i)(3-9i)=3^2+9^2=90\) and "outer and inner" the \(9i\) part will add to zero and you will get \(-3x-3x=-6x\) for a final answer of \[x^2-6x+90\]

OpenStudy (anonymous):

but if you can memorize \[x^2-2ax+a^2+b^2\] as the quadratic with zeros \(a+bi\) and \(a-bi\) then while your fellow students are struggling in their exam you can be done in like 30 seconds

OpenStudy (anonymous):

okay, now would you multiply (x^2-6x+90) and (x^2+6x-112) together?

OpenStudy (anonymous):

yeah exactly this is donkey work i would cheat

OpenStudy (anonymous):

the command "expand" will do it for this one http://www.wolframalpha.com/input/?i=expand+%28x-8%29%28x%2B14%29%28x^2-6x%2B90%29

OpenStudy (anonymous):

wow is this really your first question here?

OpenStudy (anonymous):

yes it is! and thank you for helping me!

OpenStudy (anonymous):

yw summer math, hope it is over soon for you good luck !